A mutual fund pays 3% compounded monthly. How much should I invest now so that 12 months from now I will have $3500 in the account?
Answer by Alan3354(69443)

A mutual fund pays 3% compounded monthly. How much should I invest now so that 12 months from now I will have $3500 in the account?
------------
If you mean 3% per year --> 0.0025% per month
-----
FV = PV*(1 + 0.0025)^t t in months
3500 = PV*1.0025^12
PV = 3500/(1.0025^12)
Just calculator work left